OpenStudy (lukecrayonz):

Angle between two vectors, <4,4> and <2,0>

OpenStudy (amistre64):

find dot, and magnitudes; what do you get?

OpenStudy (amistre64):

... or 45 degrees; but thats cheating

OpenStudy (lukecrayonz):

Uhh 8+0=8 for dot, and magnitude of both for what reason?

OpenStudy (amistre64):

4,4 ; 16+16 = sqrt(32) = 4sqrt(2) 2,0 ; 4+0 = sqrt(4) =2 ---- 8+0 = 8 cos(t) = 8 --------- = 1/sqrt(2) 2*4sqrt(2)

OpenStudy (lukecrayonz):

theta=acos<v1+v2>?

OpenStudy (amistre64):

it helps if you can remember the formulas ...

OpenStudy (amistre64):

a.b ----- = cos(t) |a||b|

OpenStudy (lukecrayonz):

What was I looking at then.. What formula is that I just said?! Haha

OpenStudy (amistre64):

i dunno what you posted; might be a new invention

OpenStudy (lukecrayonz):

So the answer is 45 degrees? Is there any way of using technology to find this? It's an online class, they promote online technology etc.

OpenStudy (amistre64):

ummm, isnt this technological enough? your on the internet .....

OpenStudy (lukecrayonz):

:P As in some sort of program etc

OpenStudy (amistre64):

dunno, im still trying to work out function on my abacus

OpenStudy (lukecrayonz):

Oh they provide a program to input into your calculator for this :P

OpenStudy (lukecrayonz):

I have a new question for you, you might love it.